To derive relation between various physical quantities ex. Kinetic energy depends on momentum and mass
 

Grade:11

1 Answers

Unicorn
2 years ago
Kinetic energy =1/2mv^2 , momentum(P) =mv 
KE=P^2/2m or P=root2mKE
>>h/p=lambda ,E=hc/lambda
KE= -total energy=-potential energy /2